OpenStudy (meilendurcer):

The table below shows four systems of equations: System 1 System 2 System 3 System 4 4x − 5y = 2 4x − 5y = 2 4x − 5y = 2 4x −5y= 2 3x − y = 8 3x − 8y = 4 10x − 21y = 10 10x + 3y = 15

OpenStudy (meilendurcer):

A. Which pair of systems will have the same solution? System 1 and system 2, because the second equation in system 2 is obtained by adding the first equation in system 1 to two times the second equation in system 1. B. System 2 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 2 to two times the second equation in system 2. C. System 1 and system 2, because the second equation in system 2 is obtained by adding the first equation in system 1 to three times the second equation in system 1. D. System 2 and system 3, because the second equation in system 3 is obtained by adding the first equation in system 2 to three times the second equation in system 2.

OpenStudy (meilendurcer):

@ganeshie8 can you help?

OpenStudy (perl):

did you try to use matrixes on this

OpenStudy (meilendurcer):

nope i can't solve it at all

OpenStudy (anonymous):

B is the only answer I found correct. Is the second equation of the third system equal to the first equation of system 2 plus 2 times the second equation of system 2 \[(4x - 5y = 2) + 2(3x - 8y = 4)\] \[(4x - 5y = 2) + (6x - 16y = 8)\] \[10x - 21y = 10\] - This was also the second equation of the third system

OpenStudy (meilendurcer):

@Pawtpie i was thinking it might be that one. thanks.

OpenStudy (anonymous):

No problem

OpenStudy (meilendurcer):

if you aren't busy can you help with a few more??